in an experiment to measure the value of pi,the following results are obtained for pi

3.12 incorrect 0/1 pts OTECQuestion 7 In an experiment to measure the value of pi, the following results are obtained for pi: 3.14 311320 3.06 3.08 Calculate the Average Deviation from the Mean Question 8 12 1 pts
Answer #1

N = number of values = 5

\mu = mean of the values = (3.14 + 3.11 + 3.20 + 3.06 + 3.08)/5 = 3.12

average deviation from mean is given as

ADM = ((|3.14 - 3.12|) + (|3.11 - 3.12|) + (|3.20 - 3.12|) + (|3.06 - 3.12|) + (|3.08 - 3.12|))/5

ADM = (0.02 + 0.01 + 0.08 + 0.06 + 0.04)/5

ADM = 0.042